How do I find the standard form of the equation of the line with slope -1/4 passing through the point (-6, -3)
Answer by Alan3354(69443)

How do I find the standard form of the equation of the line with slope -1/4 passing through the point (-6, -3)
================
y-y1 = m*(x-x1)
y+3 = (-1/4)*(x+6)
Convert to standard form